# Thermodynamics of Modified Chaplygin Jacobi And Modified Chaplygin Abel Gas: Stability Analysis and Observational Constraints

### v1

@@ 2.2   Deceleration Parameter
The deceleration parameter $q$ characterizes the rate at which the universe expands. A negative value of $q$ signifies that the universe is experiencing accelerated expansion, whereas a positive value indicates a phase of decelerated expansion.
The deceleration parameter is given by:
$q=\frac{1}{2}+\frac{3p}{2\rho}$ (17)
@@
$q=-\dfrac{3BK}{2\left[\frac{B}{A+1}\left(2K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)\right]^{\frac{1}{\alpha+1}}\left[\frac{B}{A+1}\left(\dfrac{\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}}{K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1}+1\right)\right]^{\frac{\alpha}{\alpha+1}}}+0.5$ (18)
@@
$\lim_{V\to 0}q$ $=0.5$ (19)
$\lim_{V\to\infty}q$ $=-\frac{3AK}{2}-\frac{3K}{2}+\frac{1}{2}$ (20)
@@
Figure 2: A graph depicting the equation of state parameter $\omega$ and the deceleration parameter $q$ as functions of volume $V$ for the Modified Chaplygin Jacobi Gas (MCJG) model.
2.3   Square of The Speed of Sound
@@
$v_{s}^{2}=\left(\frac{\partial p}{\partial\rho}\right)_{S}$
@@
$\left.-(A+1)^{2}(K-1)(\alpha+2)\left(\frac{BY}{A+1}\right)^{\frac{2\alpha+2}{\alpha+1}}\right]\left(\frac{BY}{A+1}\right)^{-\frac{\alpha+2}{\alpha+1}}$
@@
$Y=2K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1$ (22)
@@ 2.4   Thermodynamic Stability
$\left(\frac{\partial p}{\partial V}\right)_{T}=-\frac{BK\alpha\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}(A+1)}{V\left(K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)X}$ (23)
@@
$X=\left[K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right]\left[\frac{B\left(K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}{(A+1)\left(K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}\right]^{\frac{\alpha}{\alpha+1}}$ (24)
@@
$T=\dfrac{\partial U}{\partial S}$ (25)
$=\Big{(}\dfrac{\partial U}{\partial C}\Big{)}\Big{(}\dfrac{\partial C}{\partial S}\Big{)}$
The consideration of parameters $A$ and $C$ as entropy-independent constants presents a theoretical challenge in the thermodynamic framework. Under such conditions, the system would exhibit a temperature reduction to zero, regardless of the gas’s volumetric or pressure conditions. This would result in the $T=0$ isotherm simultaneously functioning as an isentropic pathway where $S=\text{const}$. Such behavior generates a fundamental inconsistency with thermodynamic principles, specifically contradicting the third law. Therefore, a comprehensive examination of the Modified Chaplygin Jacobi Gas (MCJG) thermodynamic stability necessitates the assumption that at minimum, one parameter must exhibit entropy $S$ dependence.
Now, the Internal Energy of the system is given as,
$U=V\left(\dfrac{B\left(1-\dfrac{1}{1-K-(\frac{B}{C})^{(A+1)}V^{(A+1)(\alpha+1)}}\right)}{1+A}\right)^{\frac{1}{1+\alpha}}$ (26)
@@
$C=(\tau S)^{(1+\alpha)}$ (27)
Where, we have used $U=TS$. Here, $\tau$ is the constant with the dimension of temperature.
Taking the derivative of the above equation with respect to S, we obtain
$\left(\frac{\partial C}{\partial S}\right)=(S\tau)^{\alpha+1}\frac{\alpha+1}{S}$ (28)
@@
$T=\frac{V^{(A+1)(\alpha+1)+1}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}\left[\frac{B\left(K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}\right)}{(A+1)\left(K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}-1\right)}\right]^{\frac{1}{\alpha+1}}(A+1)}{S\left[K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}\right]\left[K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}-1\right]}$ (29)
@@
$C_{v}=T\left(\dfrac{\partial S}{\partial T}\right)_{V}$ (30)
@@
$C_{V}=\frac{SV^{\beta+1}\Gamma^{A+1}(A+1)(K+\zeta)(K+\zeta-1)}{V^{\beta+1}\Gamma^{A+1}(K+\zeta)[-\xi(K+\zeta-1)+V^{\beta}\Gamma^{A+1}(A+1)^{2}(2\alpha+3+\frac{1}{K+\zeta})]}$ (31)
@@
$\beta=(A+1)(\alpha+1)$
$\Gamma=(B(S\tau)^{-\alpha-1})^{A+1}$
$\zeta=V^{\beta}\Gamma$
$\xi=A+\beta+1=(A+1)(2\alpha+3)$
The positive value of specific heat at constant volume $C_{V}$ is shown in the fig. 4(a). This, together with the previous conditions of $\left(\dfrac{\partial p}{\partial V}\right)_{T}<0$ ensure that MCJG is thermodynamically stable.
Again, the temperature is given as,
$T=\dfrac{p+\rho}{S}$ (32)
@@
$T=-\frac{\dfrac{BK}{\left[B\left(\frac{(V_{0}/V)^{(A+1)(\alpha+1)}}{K(V_{0}/V)^{(A+1)(\alpha+1)}+1}+1\right)/(A+1)\right]^{\frac{\alpha}{\alpha+1}}}-\left(\frac{B(2K(V_{0}/V)^{(A+1)(\alpha+1)}+1)}{A+1}\right)^{\frac{1}{\alpha+1}}}{S}$ (33)
@@ 2.5   Reduced Parameters
$p(T)=-\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}$ (35)
@@
$\left(\frac{C}{B}\right)^{(1+A)}=\left[\left(\frac{A+1}{B}\right)\rho_{i}^{1+\alpha}-1\right]\left(\frac{1}{2K}\right)V_{i}^{(1+A)(1+\alpha)}$ (36)
@@
$\rho=\rho_{i}\left[\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{(1+A)(1+\alpha)}\right]^{\frac{1}{1+\alpha}}$ (37)
$p=-BK\left[\rho_{i}\left(\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{(1+A)(1+\alpha)}\right)^{\frac{1}{1+\alpha}}\right]^{-\alpha}$ (38)
@@
$\epsilon$ $=$ $\dfrac{\rho}{\rho_{i}}$ (39)
$v$ $=$ $\dfrac{V}{V_{i}}$ (40)
$P$ $=$ $\dfrac{p}{\left(\frac{B}{A+1}\right)^{-\frac{\alpha}{1+\alpha}}}$ (41)
$\gamma$ $=$ $\dfrac{B}{(A+1)\rho_{i}^{1+\alpha}}$ (42)
$t$ $=$ $\dfrac{T}{T_{i}}$ (43)
$t^{*}$ $=$ $\dfrac{T^{*}}{T_{i}}$ (44)
@@
$\epsilon=\left[\gamma+\dfrac{1-\gamma}{v^{(1+A)(1+\alpha)}}\right]^{\frac{1}{1+\alpha}}$ (45)
$P=-BK\gamma^{\frac{\alpha}{1+\alpha}}\cdot\left[\gamma+\dfrac{1-\gamma}{v^{(1+A)(1+\alpha)}}\right]^{-\frac{\alpha}{1+\alpha}}$ (46)
$p(T)=-\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{t}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}$ (47)
@@
$\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{1}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}=BK\gamma^{\dfrac{\alpha}{\alpha+1}}$ (48)
@@
$t^{*}=\left(\dfrac{1}{BK}\right)^{\frac{1}{1+\alpha}}\left[\dfrac{1}{1-\gamma}\right]^{\frac{\alpha}{\alpha+1}}$ (49)
@@
$\gamma=1-(BK)^{-\frac{1}{\alpha}}\left(t^{*}\right)^{-\frac{\alpha+1}{\alpha}}$ (50)
In our analysis, we consider two critical temperature points in the cosmological timeline: the theoretical peak temperature of the Modified Chaplygin Jacobi Gas (MCJG), denoted as $T^{*}$, which is estimated at $10^{32}$ units, coinciding with temperature conditions during the Planck epoch. This is contrasted with the contemporary cosmic microwave background temperature, measured at $T_{i}=2.7$ units. From these boundary conditions, we can proceed with our calculations.
$\gamma=1-(BK)^{-\frac{1}{\alpha}}\left(10^{32}\right)^{-\frac{\alpha+1}{\alpha}}\approx 1$ (51)
Consequently, in this scenario, the dark energy density of the universe permeated with Modified Chaplygin Jacobi gas is nearly equal to its limiting value, given by $\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}$. The internal energy of the system becomes,
$U=\dfrac{(BK)^{\frac{1}{\alpha}}V}{\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}$ (52)
@@
$\dfrac{\partial U}{\partial V}=T\dfrac{\partial p}{\partial T}-p$ (53)
@@
$\dfrac{(BK)^{\frac{1}{\alpha}}}{\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}=\dfrac{BK}{\rho^{\alpha}}\left[\frac{T}{p}\frac{dp}{dT}-1\right]$ (54)
@@
3   Thermodynamic Analysis of Modified Chaplygin Abel Gas
By substituting the hyperbolic function with the Abel elliptic function, we can derive the pressure-density relationship of the Modified Chaplygin Abel Gas as[44],
$p=\left[(e^{2}+2c^{2})(1+A)-1\right]\rho-\frac{c^{2}B}{\rho^{\alpha}}-\frac{(c^{2}+e^{2})(1+A)^{2}\rho^{2+\alpha}}{B}$ (55)
Here, $A$, $B$ are positive constant parameters, $0\leq\alpha\leq 1$ and $c$, $e$ are real numbers. This is the pressure-density relation of “Modified Chaplygin Abel Gas (MCAG)”.
Now, using equation (3) and (4) we get the Internal Energy as follows:
$U=V\left[\dfrac{Bc^{2}}{(1+A)\left(c^{2}+e^{2}\left(1+\dfrac{1}{-1+c^{2}\exp((1+A)(1+\alpha)Be^{2}d)V^{(1+A)(1+\alpha)e^{2}}}\right)\right)}\right]^{\frac{1}{1+\alpha}}$ (56)

### v2

@@ 2.2   Deceleration Parameter
The deceleration parameter $q$ characterizes the rate at which the universe expands. A negative value of $q$ signifies that the universe is experiencing accelerated expansion, whereas a positive value indicates a phase of decelerated expansion. The deceleration parameter is given by:
$q=\frac{1}{2}+\frac{3p}{2\rho}\quad.$ (14)
@@
$q=-\dfrac{3BK}{2\left[\frac{B}{A+1}\left(2K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)\right]^{\frac{1}{\alpha+1}}\left[\frac{B}{A+1}\left(\dfrac{\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}}{K\left(\dfrac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1}+1\right)\right]^{\frac{\alpha}{\alpha+1}}}+0.5\quad.$ (15)
@@
$\lim_{V\to 0}q$ $=0.5,$ (16)
$\lim_{V\to\infty}q$ $=-\frac{3AK}{2}-\frac{3K}{2}+\frac{1}{2}\quad.$ (17)
@@
Figure 2: A graph depicting the equation of state parameter $\omega$ and the deceleration parameter $q$ as functions of volume $V$ for the Modified Chaplygin-Jacobi Gas model.
2.3   Square Speed of Sound
@@
$v_{s}^{2}=\left(\frac{\partial p}{\partial\rho}\right)_{S}.$
@@
$\left.-(A+1)^{2}(K-1)(\alpha+2)\left(\frac{BY}{A+1}\right)^{\frac{2\alpha+2}{\alpha+1}}\right]\left(\frac{BY}{A+1}\right)^{-\frac{\alpha+2}{\alpha+1}}\quad,$
@@
$Y=2K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\quad.$ (19)
@@ 2.4   Thermodynamic Stability
$\left(\frac{\partial p}{\partial V}\right)_{T}=-\frac{BK\alpha\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}(A+1)}{V\left(K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)X}\quad,$ (20)
@@
$X=\left[K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right]\left[\frac{B\left(K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}{(A+1)\left(K\left(\frac{V_{0}}{V}\right)^{(A+1)(\alpha+1)}+1\right)}\right]^{\frac{\alpha}{\alpha+1}}.$ (21)
@@
$T=\dfrac{\partial U}{\partial S}=\Big{(}\dfrac{\partial U}{\partial C}\Big{)}\Big{(}\dfrac{\partial C}{\partial S}\Big{)}.$ (22)
The consideration of parameters $A$ and $C$ as entropy-independent constants presents a theoretical challenge in the thermodynamic framework. Under such conditions, the system would exhibit a temperature reduction to zero, regardless of the gas’s volumetric or pressure conditions. This would result in the $T=0$ isotherm simultaneously functioning as an isentropic pathway where $S=\text{const}$. Such behavior generates a fundamental inconsistency with thermodynamic principles, specifically contradicting the third law. Therefore, a comprehensive examination of the MCJG thermodynamic stability necessitates the assumption that at minimum, one parameter must exhibit entropy $S$ dependence. Now, the Internal Energy of the system is given as,
$U=V\left(\dfrac{B\left(1-\dfrac{1}{1-K-(\frac{B}{C})^{(A+1)}V^{(A+1)(\alpha+1)}}\right)}{1+A}\right)^{\frac{1}{1+\alpha}}.$ (23)
@@
$C=(\tau S)^{(1+\alpha)},$ (24)
where, we have used $U=TS$. Here, $\tau$ is the constant with the dimension of temperature. Taking the derivative of the above equation with respect to S, we obtain
$\left(\frac{\partial C}{\partial S}\right)=(S\tau)^{\alpha+1}\frac{\alpha+1}{S}\quad.$ (25)
@@
$T=\frac{V^{(A+1)(\alpha+1)+1}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}\left[\frac{B\left(K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}\right)}{(A+1)\left(K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}-1\right)}\right]^{\frac{1}{\alpha+1}}(A+1)}{S\left[K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}\right]\left[K+V^{(A+1)(\alpha+1)}\left(B(S\tau)^{-\alpha-1}\right)^{A+1}-1\right]}\quad.$ (26)
@@
$C_{v}=T\left(\dfrac{\partial S}{\partial T}\right)_{V}.$ (27)
@@
$C_{V}=\frac{SV^{\beta+1}\Gamma^{A+1}(A+1)(K+\zeta)(K+\zeta-1)}{V^{\beta+1}\Gamma^{A+1}(K+\zeta)[-\xi(K+\zeta-1)+V^{\beta}\Gamma^{A+1}(A+1)^{2}(2\alpha+3+\frac{1}{K+\zeta})]}\quad,$ (28)
@@
$\beta=(A+1)(\alpha+1),\quad\Gamma=(B(S\tau)^{-\alpha-1})^{A+1},\quad\zeta=V^{\beta}\Gamma,\quad\xi=A+\beta+1=(A+1)(2\alpha+3).$
The positive value of specific heat at constant volume $C_{V}$ is shown in the fig. 4(a). This, together with the previous conditions of $\left(\dfrac{\partial p}{\partial V}\right)_{T}<0$ ensure that MCJG is thermodynamically stable. Again, the temperature is given as,
$T=\dfrac{p+\rho}{S}\quad,$ (29)
@@
$T=-\frac{\dfrac{BK}{\left[B\left(\frac{(V_{0}/V)^{(A+1)(\alpha+1)}}{K(V_{0}/V)^{(A+1)(\alpha+1)}+1}+1\right)/(A+1)\right]^{\frac{\alpha}{\alpha+1}}}-\left(\frac{B(2K(V_{0}/V)^{(A+1)(\alpha+1)}+1)}{A+1}\right)^{\frac{1}{\alpha+1}}}{S}\quad.$ (30)
@@ 2.5   Reduced Parameters
$p(T)=-\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}.$ (32)
@@
$\left(\frac{C}{B}\right)^{(1+A)}=\left[\left(\frac{A+1}{B}\right)\rho_{i}^{1+\alpha}-1\right]\left(\frac{1}{2K}\right)V_{i}^{(1+A)(1+\alpha)}.$ (33)
@@
$\rho=\rho_{i}\left[\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{(1+A)(1+\alpha)}\right]^{\frac{1}{1+\alpha}}\quad,$ (34)
$p=-BK\left[\rho_{i}\left(\frac{B}{(A+1)\rho_{i}^{1+\alpha}}+\left(1-\frac{B}{(A+1)\rho_{i}^{1+\alpha}}\right)\left(\frac{V_{i}}{V}\right)^{(1+A)(1+\alpha)}\right)^{\frac{1}{1+\alpha}}\right]^{-\alpha}\quad.$ (35)
@@
$\epsilon=\dfrac{\rho}{\rho_{i}},\quad v=\dfrac{V}{V_{i}},\quad P=\dfrac{p}{\left(\frac{B}{A+1}\right)^{-\frac{\alpha}{1+\alpha}}},\quad\gamma=\dfrac{B}{(A+1)\rho_{i}^{1+\alpha}},\quad t=\dfrac{T}{T_{i}},\quad t^{*}=\dfrac{T^{*}}{T_{i}}.$ (36)
@@
$\epsilon=\left[\gamma+\dfrac{1-\gamma}{v^{(1+A)(1+\alpha)}}\right]^{\frac{1}{1+\alpha}},$ (37)
$P=-BK\gamma^{\frac{\alpha}{1+\alpha}}\cdot\left[\gamma+\dfrac{1-\gamma}{v^{(1+A)(1+\alpha)}}\right]^{-\frac{\alpha}{1+\alpha}},$ (38)
$p(T)=-\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{t}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}.$ (39)
@@
$\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{1}{t^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{\alpha}{\alpha+1}}=BK\gamma^{\frac{\alpha}{\alpha+1}}.$ (40)
@@
$t^{*}=\left(\dfrac{1}{BK}\right)^{\frac{1}{1+\alpha}}\left[\dfrac{1}{1-\gamma}\right]^{\frac{\alpha}{\alpha+1}},$ (41)
@@
$\gamma=1-(BK)^{-\frac{1}{\alpha}}\left(t^{*}\right)^{-\frac{\alpha+1}{\alpha}}.$ (42)
In our analysis, we consider two critical temperature points in the cosmological timeline: the theoretical peak temperature of the Modified Chaplygin-Jacobi Gas, denoted as $T^{*}$, which is estimated at $10^{32}$ units, coinciding with temperature conditions during the Planck epoch. This is contrasted with the contemporary cosmic microwave background temperature, measured at $T_{i}=2.7$ units. From these boundary conditions, we can proceed with our calculations.
$\gamma=1-(BK)^{-\frac{1}{\alpha}}\left(10^{32}\right)^{-\frac{\alpha+1}{\alpha}}\approx 1\quad.$ (43)
Consequently, in this scenario, the dark energy density of the universe permeated with Modified Chaplygin-Jacobi gas is nearly equal to its limiting value, given by $\left(\dfrac{B}{A+1}\right)^{\frac{1}{\alpha+1}}$. The internal energy of the system becomes,
$U=\dfrac{(BK)^{\frac{1}{\alpha}}V}{\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}\quad.$ (44)
@@
$\dfrac{\partial U}{\partial V}=T\dfrac{\partial p}{\partial T}-p\quad,$ (45)
@@
$\dfrac{(BK)^{\frac{1}{\alpha}}}{\left[(BK)^{\frac{1}{\alpha}}-\left(\frac{T}{T^{*}}\right)^{\frac{\alpha+1}{\alpha}}\right]^{\frac{1}{\alpha+1}}}=\dfrac{BK}{\rho^{\alpha}}\left[\frac{T}{p}\frac{dp}{dT}-1\right]\quad.$ (46)
@@
3   Thermodynamic Analysis of Modified Chaplygin-Abel Gas
By substituting the hyperbolic function with the Abel elliptic function, we can derive the pressure-density relationship of the Modified Chaplygin-Abel gas as[44],
$p=\left[(e^{2}+2c^{2})(1+A)-1\right]\rho-\frac{c^{2}B}{\rho^{\alpha}}-\frac{(c^{2}+e^{2})(1+A)^{2}\rho^{2+\alpha}}{B}\quad.$ (47)
Here, $A$, $B$ are positive constant parameters, $0\leq\alpha\leq 1$ and $c$, $e$ are real numbers. This is the pressure-density relation of “Modified Chaplygin-Abel Gas”. The physical significances of $A$, $B$ and $\alpha$ parameters are detailed in Sect. 2. Though the elliptic parameters $e$, $c$ have no distinguished physical significance but they greatly affect the pressure-density relation of the modeled dark energy and henceforth have a deep impact on the evolution dynamics of the universe. The specific choice of these parameters can reduce the eq. (47) to another Chaplygin gas equation of state which we will see in the next paragraph. The observational constraints of their value is explicitly studied in [44]. Now, using equation (2) and (3) we get the Internal Energy as follows:
$U=V\left[\dfrac{Bc^{2}}{(1+A)\left(c^{2}+e^{2}\left(1+\dfrac{1}{-1+c^{2}\exp((1+A)(1+\alpha)Be^{2}d)V^{(1+A)(1+\alpha)e^{2}}}\right)\right)}\right]^{\frac{1}{1+\alpha}}.$ (48)